OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN PLEASE HELP! The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank. 1 over 3π (x - 7)2 (x - 3) What does the factor π(x - 7)2 (x - 3) represent?

OpenStudy (raden):

volume of a cone = 1/3 * pi * base area * height

OpenStudy (raden):

it means, radius = (x-7) height = (x-3)